# Enhanced Athermal Phonon Responsivity in a Kinetic Inductance Detector with Integrated Phonon Collectors

### v1

@@ Abstract
Experiments searching for direct light dark matter interaction Abdelhameed et al. (2019); Armengaud et al. (2016); Agnese et al. (2018) or coherent elastic neutrino-nucleus scattering Strauss et al. (2017); Billard et al. (2017); Augier et al. (2023); Agnolet et al. (2017); Akimov et al. (2017) have to address two main technological challenges: the need for high sensitivity to nuclear recoils and the scaling of the instrumented target mass.
The small nuclear recoil energies require detectors with energy thresholds as low as possible. For this reason, cryogenic detectors such as Transition Edge Sensors (TESs) Abdelhameed et al. (2019); Strauss et al. (2017); Alkhatib et al. (2021) or Neutron Transmutation Doped (NTD) Germanium thermistors Armengaud et al. (2016); Augier et al. (2023) are adopted as phonon sensors coupled to target crystals, in order to detect sub-keV nuclear recoil processes. Energy thresholds of the order of $1-10\,\mathrm{eV}$ have already been achieved Anthony-Petersen et al. (2025); Strauss et al. (2017); Alkhatib et al. (2021).
@@
At the same time, the development of mass-scalable instruments is essential for achieving detectors with improved sensitivity to increasingly smaller dark-matter interaction cross sections, assuming a Weakly Interacting Massive Particle (WIMP) model. Experiments with target mass of tens of grams have been built Abdelhameed et al. (2019); Strauss et al. (2017); Armengaud et al. (2019), but scaling-up to the kilogram still remains an experimental challenge.
@@
The BULLKID project developed a particle detector designed for tackling the mass scaling problem Cruciani et al. (2022). It consists of a monolithic array of 60 silicon dices, each one with mass of 0.34 g and volume $5.4\times 5.4\times 5\,\mathrm{mm^{3}}$, carved from a wafer with a diameter of $7.6\,\mathrm{cm}$ and 5 mm thickness. The dices are sensed with 60 KIDs acting as phonon-mediated particle detectors. These are well suited for use in large arrays due to their intrinsic multiplexing capability Day et al. (2003) and therefore can be adopted to design mass-scalable instruments. BULLKID achieved an energy resolution of $27\,\mathrm{eV}$ — which led to set an analysis threshold of around $160~\mathrm{eV}$ Delicato et al. (2024) — that is still far from the current state of art of the other cryogenic detectors. For this reason, it is important to design and develop new sensors aimed at lowering the energy threshold, to the greatest extent possible. In this letter, we investigate a KID design aimed at achieving this goal.
@@
KIDs, which were originally developed for the detection of optical Mazin et al. (2019) and millimeter-wavelength Adam, R. et al. (2018) light, have found application also as phonon-mediated particle detectors Cruciani et al. (2022); Swenson et al. (2010); Cruciani et al. (2012); Golwala et al. (2008). A KID is a superconducting LC resonator Zmuidzinas (2012) in which the total inductance $L$ is the sum of two contributions: the magnetic inductance $L_{g}$, which is related to the geometry of the superconductor, and the kinetic inductance $L_{k}$, which is associated to the Cooper Pairs (CPs) motion inside the metal and depends on their number density. When energy $E$ is released in the substrate onto which the phonon-mediated KID is deposited, athermal phonons are generated and absorbed by the KID itself. If the phonon energy $E_{\mathrm{ph}}$ satisfies $E_{\mathrm{ph}}>2\Delta_{0}$, where $\Delta_{0}\simeq 1.76k_{B}T_{c}$ is the superconducting gap of the metal, $k_{B}$ the Boltzmann constant and $T_{c}$ the critical temperature  Zmuidzinas (2012), CPs in the superconductor are broken and Quasi-Particles (QPs) are generated. This process reduces the CP number density, which in turn increases the kinetic inductance $L_{k}$, leading to a resonant frequency shift, $\Delta f_{r}$, and to a reduction of the resonance depth Zmuidzinas (2012); Mazin (2005); Khalil et al. (2012); Swenson et al. (2013); Vignati et al. (2021).
@@
The change of the CP density is detected by monitoring the phase and magnitude modulation of the complex single pole transmission function $S_{21}(\Delta f_{r})$ of the resonator Delicato et al. (2025):
$S_{21}(\Delta f_{r})=1-\frac{Q}{Q_{c}}\frac{1}{1+2jy}$ (1)
@@
where $Q=(Q_{c}^{-1}+Q_{i}^{-1})^{-1}$ is the total quality factor of the resonator, which combines the coupling quality factor $Q_{c}$ and the internal quality factor $Q_{i}$, that accounts for internal losses due to imperfections and thermal QPs population within the sensor. Finally, the parameter $y=Q\cdot\Delta f_{r}/f_{r}$ is the detuning of the readout frequency relative to the resonance width Zmuidzinas (2012); Mazin (2005); Vignati et al. (2021) Usually, the phase readout $\varphi$ provides a better signal-to-noise ratio Zmuidzinas (2012); therefore, this will be adopted in this work.
The signal responsivity $r$ of a KID in terms of $\varphi$ is Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008); Delicato et al. (2025):
$r=\frac{d\varphi}{dE}=\frac{\eta}{V_{\mathrm{KID}}}\frac{\alpha S_{2}(f_{r},T)Q}{N_{0}\Delta_{0}^{2}}\frac{1}{1+4y^{2}}$ (2)
@@
where $\eta$ is the energy to pair-breaking conversion efficiency, $V_{\mathrm{KID}}$ is the active volume of the sensor, $\alpha=L_{k}/(L_{k}+L_{g})$ is the fraction of kinetic inductance, $S_{2}(f_{r},T)$ is a dimension-less function of order 1 coming from the Mattis-Bardeen theory and $N_{0}$ is the the single-spin density of states. We emphasize that in Eq. 2 the term $\Delta_{0}$ appears quadratically as the product of two identical gap factors, $\Delta_{0}\cdot\Delta_{0}$: the first $\Delta_{0}$ arises from the number of CPs broken by a given energy deposition $E$, which scales as $E/\Delta_{0}$, while the second $\Delta_{0}$ originates from the single–spin volume density, $N_{0}\Delta_{0}$, within the sensor Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008).
@@
If the phonons absorption is not saturated and losses due to supports or interfaces do not dominate, the efficiency $\eta$ scales almost linearly with the phonon collection volume $V_{\mathrm{ph}}$ Martinez et al. (2019); Fyhrie et al. (2016); Santis (2022):
@@
In a standard phonon-mediated KID, the volume $V_{\mathrm{ph}}$ coincides with that of the active sensor $V_{\mathrm{KID}}$, since the phonons generated in the substrate are directly absorbed by the inductor. Consequently, from Eq. 2 it follows that the responsivity $r$ of the KID weakly depends on $V_{\mathrm{KID}}$, thus tuning the active sensor volume alone is not expected to lead to a significant improvement in the detector responsivity. In contrast, the separation of these two volumes would allow to design a device with a large $V_{\mathrm{ph}}$ and a small $V_{\mathrm{KID}}$ at the same time, which would enhance the detector responsivity.
@@
We developed and tested a phonon-mediated KID with separate phonon collectors Santis (2022); Mazin et al. (2006); Pesce et al. (2026) that act also as QPs funnels (see Fig. 1). Henceforth, we will refer to this device as the FunKID and to the phonon collectors as the funnels. The resonator is composed of a $77\,\mathrm{nm}$ trilayer meander of Aluminum-Titanium-Aluminum (AlTiAl), with $14\,\mathrm{nm}-33\,\mathrm{nm}-30\,\mathrm{nm}$ thicknesses, respectively Cardani et al. (2018). The inductive part of the KID is composed of seven fingers of around $3600\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}$, evenly spaced by $215\,\mathrm{\mu m}$ (Fig. 1(a)). The funnels are Aluminum structures with $210\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}\times 100\,\mathrm{nm}$ volume intersecting the inductor, evenly spaced by $10\,\mathrm{\mu m}$ (Fig. 1(b)). The total phonon collection volume, which coincides with the total volume of the funnels, is $V_{\mathrm{ph}}\simeq 5.2\,\mathrm{mm}^{2}\times 100\,\mathrm{nm}$, while the active volume of the sensor is $V_{\mathrm{KID}}\simeq 0.55\,\mathrm{mm}^{2}\times 77\,\mathrm{nm}$, giving a volume ratio of $V_{\mathrm{ph}}/V_{\mathrm{KID}}\simeq 12$.
@@
Figure 1: (a) Layout of the FunKID. The funnels in purple intersect the inductive meander indicated with L. The capacity C is adjusted by means of the first two capacitors to tune the resonant frequency. (b) Detail of the funnels. The measures are in $\mathrm{\mu m}$.
@@
Due to proximity effects Cardani et al. (2018); Catalano, A. et al. (2015); Zhao et al. (2017); Brammertz et al. (2001) among the two Aluminum layers ($T_{c}\simeq 1.2\,\mathrm{K}$ Zhao et al. (2017)) and the central Titanium layer ($T_{c}\simeq 0.4\,\mathrm{K}$ Zhao et al. (2017)), the AlTiAl sensor has an intermediate critical temperature $T_{c}$, which is lower than that of the funnels, made of Aluminum. Generally, the AlTiAl critical temperature depends on the boundary resistivities conditions between the layers, and on their thicknesses Zhao et al. (2018). For the same design thicknesses, the CALDER project obtained a value of $T_{c}\simeq 0.81\,\mathrm{K}$ Cardani et al. (2018, 2021), which is expected also for this resonator, provided that the boundary resistance conditions are sufficiently similar.
@@
In this design, schematically reported Fig. 2(a), the athermal phonons produced in the substrate are mostly absorbed by the funnels, where they can break CP generating QPs, which, through diffusion, reach the sensor. Due to the energy gap difference between funnels with gap $\Delta_{\mathrm{Al}}$ and sensor with gap $\Delta_{\mathrm{AlTiAl}}$, QPs with energy $\Delta_{\mathrm{Al}}$ relax down to $\Delta_{\mathrm{AlTiAl}}$ and are trapped within the sensor, preventing their recombination in the funnels Mazin et al. (2006); Riwar and Catelani (2019).
@@
The length of the funnels was chosen to be smaller than the characteristic QP diffusion length in aluminum, $\lambda_{r}=\sqrt{D\tau_{r}}$ Nussbaumer et al. (2000), where $D$ is the QP diffusion coefficient and $\tau_{r}$ the QP recombination time. We adopted a diffusion coefficient of $D=22.5\,\mathrm{cm^{2}/s}$ Hsieh and Levine (1968), a conservative choice compared to higher values reported in the literature Zgirski et al. (2020). The recombination time $\tau_{r}$ depends on both film thickness and temperature, and we adopted $\tau_{r}\simeq 1\,\mathrm{ms}$ de Visser et al. (2011); Barends et al. (2008), yielding a diffusion length of $\lambda_{r}\simeq 1.5\,\mathrm{mm}$, much higher than half the length of the funnels of $100\,\mathrm{\mu m}$. We can also define a characteristic time scale that quantifies the maximum diffusion time of QPs inside the funnels, $\tau_{\mathrm{diff}}\sim l^{2}/D=4\,\mathrm{\mu s}$, where $l=100\,\mathrm{\mu m}$ is half of the funnel length and corresponds to the maximum distance a QP travels before reaching the sensor. This sets the minimum timescale over which the sensor needs to integrate the signal.
@@
Figure 2: (a) Detection scheme of the FunKID. Athermal phonons (white dots) are generated by an energy deposition in the substrate. A fraction of these phonons is absorbed by the funnels, which have a higher superconducting gap $\Delta_{\mathrm{Al}}$. There, CPs (red paired dots) are broken, producing QPs (single red dots). Finally, the QPs diffuse into the funnel and are absorbed by the KID, which features a lower gap $\Delta_{\mathrm{AlTiAl}}$, where they recombine. (b) Current density simulated with SONNET. Most of the current flows into the KID, while a small leakage is observed. The length reported is measured in $\mathrm{\mu m}$.
@@
We estimated the gain in responsivity of the FunKID relative to an identical standard KID without funnels, $r_{\mathrm{FunKID}}/r_{\mathrm{KID}}$, using the Eq. 2:
$\frac{r_{\mathrm{FunKID}}}{r_{\mathrm{KID}}}\simeq\frac{V_{\mathrm{ph}}}{V_{\mathrm{KID}}}\frac{\alpha_{\mathrm{FunKID}}}{\alpha_{\mathrm{KID}}}\frac{Q_{\mathrm{FunKID}}}{Q_{\mathrm{KID}}}\frac{\Delta_{\mathrm{Al}}}{\Delta_{\mathrm{AlTiAl}}}$ (4)
where we assumed that $N_{0}$ and $S_{2}(f_{r},T)$ were approximately equal among the two devices and that their response was linear, thus $1+4y^{2}\simeq 1$. Moreover, we considered only the geometric dependence of $\eta$ for both the KIDs according to Eq. 3 and that the phonon collection volume of the standard KID coincides with its active volume $V_{\mathrm{KID}}$. In order to evaluate the best gain in responsivity, we assumed same quality factors $Q$ and kinetic inductance fractions $\alpha$ for both the devices and the critical temperatures of Aluminum and AlTiAl of $1.2\,\mathrm{K}$ and $0.81\,\mathrm{K}$, respectively. Under these conditions, the expected gain amounts to about 8.
The test device was fabricated at the PTA facility clean room in Grenoble Plateforme technologique amont using a $500\,\mathrm{\mu m}$ thick silicon tile. On the same substrate, a standard phonon-mediated KID was produced along with the FunKID (Fig. 3). The KID sensors were etched from a trilayer of Aluminum ($30\,\mathrm{nm}$), Titanium ($14\,\mathrm{nm}$) and Aluminum ($13\,\mathrm{nm}$) following a standard lithographic technique. After the fabrication of the KIDs, photo-resist was used to coat the surface of the wafer, and gaps in the shape of the funnels were patterned on it. The funnels were then realized by selectively depositing $100\,\mathrm{nm}$ of Aluminum in the patterned gaps, following a lift-off process. The two resonators are identical except for the first two capacitors length: in the FunKID these were designed $105\,\mathrm{\mu m}$ longer to lower its resonant frequency by approximately $20\,\mathrm{MHz}$ with respect to the KID. Finally, we highlight that the feed-line is fabricated from the same AlTiAl trilayer used for the meanders of the two KIDs.
@@
The tile was mounted into a copper holder by means of Teflon supports and wire bonded to $50\,\mathrm{\Omega}$ launchers. The device was operated in a dilution $^{3}\mathrm{He}/^{4}\mathrm{He}$ cryostat, at a base temperature $T_{0}\simeq 35\,\mathrm{mK}$, much lower than the AlTiAl critical temperature. The copper holder was closed and covered by aluminum foil to mitigate environmental radiation. Finally, a mu-metal shield was placed outside the cryostat to reduce magnetic-field effects Flanigan et al. (2016).
To calibrate the detectors and characterize their pulses, three optical fibers were placed facing the side of the silicon tile opposite to the KIDs: two positioned to face the center of each resonator, and one located midway between them. The two lateral fibers were used to identify the two pixels, while the central one was used to calibrate the KIDs from the same position for comparison. The fibers were pulsed with $400\,\mathrm{nm}$ photons bursts generated at room temperature by controlled a LED lamp Del Castello (2024).
The readout tones were generated by the NIXA electronic system Bourrion et al. (2011, 2013), a superheterodyne radio-frequency detection system capable to transmit and receive tones at the KIDs resonant frequencies. The transmitted waves undergo an attenuation chain of $55\,\mathrm{dB}$ before the chip, are transmitted past the resonators, then are amplified by $30\,\mathrm{dB}$ by a Low Noise Amplifier (model: CITLF3 Cosmic Microwave Technology, Inc. (2025)) operated at $4\,\mathrm{K}$ and finally received back by NIXA.
The resonances of the two KIDs are shown in Fig. 4. We fitted both the amplitude of the transmission function $|S_{21}|$ and its real and imaginary components in order to extract the resonant frequencies of the two KIDs as well as the quality factors $Q_{c}$, $Q_{i}$, and $Q$. We also evaluated the resonator ring times $\tau_{\mathrm{ring}}=Q/\pi f_{r}$ of the two KIDs. The results are reported in Tab. 1.
@@
Table 1: Results of the resonant frequency $f_{r}$, quality factors $Q$, $Q_{i}$ and $Q_{c}$ and ring times $\tau_{\mathrm{ring}}$ of the two resonators. The uncertainty of the resonant frequency is negligible, while the relative uncertainty on the internal quality factor $Q_{i}$ was estimated to around $10\,\%$.
A shift of around $40\,\mathrm{MHz}$ of the measured FunKID resonant frequency with respect to the simulation was observed, indicating that the overlap regions between the funnels and the resonator might have not been correctly modeled in the simulations. Conversely, the obtained KID resonant frequency is in agreement with simulation, assuming a kinetic inductance $L_{k}=1.4\,\mathrm{pH/sq}$, which is the same of Ref. Cardani et al. (2018). The coupling quality factors $Q_{c}\sim 20\,\mathrm{k}$ are in line with the values simulated with SONNET, for both resonators. In addition, the internal quality factor of the FunKID is comparable to the coupling quality factor, which currently limits this design. Conversely, the internal quality of the KID is higher than the FunKID and in line with the values obtained in Ref. Cardani et al. (2018), indicating that the reduced $Q_{i}$ of the FunKID is set by the funnels and not by fabrication processes of the trilayer AlTiAl film.
For each resonator, we evaluated our best estimate of kinetic inductance fraction $\alpha$, by combining the measured resonant frequencies $f_{r}$ with the resonant frequencies $f_{0}$ obtained from SONNET simulations performed with a perfect conductor (i.e., with $L_{k}=0$), through the formula:

### v2

@@ Abstract
Experiments searching for direct light dark matter interaction Abdelhameed et al. (2019); Armengaud et al. (2016); Agnese et al. (2018) or coherent elastic neutrino-nucleus scattering Strauss et al. (2017); Billard et al. (2017); Augier et al. (2023); Agnolet et al. (2017); Akimov et al. (2017) address two main technological challenges: the need for high sensitivity to nuclear recoils and the scaling of the instrumented target mass.
@@
The small nuclear recoil energies require detectors with energy thresholds as low as possible. For this reason, cryogenic detectors Abdelhameed et al. (2019); Armengaud et al. (2016); Strauss et al. (2017); Augier et al. (2023); Alkhatib et al. (2021) are adopted as phonon sensors coupled to target crystals, while energy thresholds of the order of 1-10 eV have already been achieved Anthony-Petersen et al. (2025); Strauss et al. (2017); Alkhatib et al. (2021). At the same time, the development of mass-scalable instruments is essential for achieving detectors with improved sensitivity to increasingly smaller dark-matter interaction cross sections, assuming a Weakly Interacting Massive Particle (WIMP) model Roszkowski et al. (2018). Experiments with target mass of tens of grams have been built Abdelhameed et al. (2019); Strauss et al. (2017); Armengaud et al. (2019), but scaling-up to the kilogram still remains an experimental challenge.
@@
The BULLKID project developed a particle detector designed for tackling the mass scaling problem Cruciani et al. (2022). It consists of a monolithic array of 60 silicon dices, each one with mass of 0.34 g and volume $5.4\times 5.4\times 5\,\mathrm{mm^{3}}$, carved from a wafer with a diameter of 7.6 cm and 5 mm thickness. The dices are sensed with 60 KIDs acting as phonon-mediated particle detectors. The best energy resolution achieved amounts to 27 eV — which led to set an analysis threshold of around 160 eV Delicato et al. (2024) — which is still far from the current state of art of the other cryogenic detectors. For this reason, it is important to develop new sensors aimed at lowering the energy threshold, to the greatest extent possible. In this letter, we investigate a KID design aimed at achieving this goal.
@@
KIDs, which were originally developed for the detection of optical Mazin et al. (2019) and millimeter-wavelength Adam, R. et al. (2018) light, have found application also as phonon-mediated particle detectors Cruciani et al. (2022); Swenson et al. (2010); Cruciani et al. (2012); Golwala et al. (2008). A KID is a superconducting LC resonator Zmuidzinas (2012) in which the total inductance $L$ is the sum of the magnetic inductance $L_{g}$ and the kinetic inductance $L_{k}$. When energy $E$ is released in the substrate of the KID, athermal phonons are generated and absorbed by the KID, where they can break Cooper Pairs (CPs) releasing Quasi-Particles (QPs). This process reduces the CP number density, which in turn increases the kinetic inductance $L_{k}$, leading to a resonant frequency shift, $\Delta f_{r}$, and to a reduction of the resonance depth Zmuidzinas (2012); Mazin (2005); Khalil et al. (2012); Swenson et al. (2013); Vignati et al. (2021).
@@
The change of the CP density is detected by monitoring the phase and magnitude modulation of the complex single pole transmission function $S_{21}(y)$ of the resonator Delicato et al. (2025):
$S_{21}(y)=1-\frac{Q}{Q_{c}}\frac{1}{1+2jy}$ (1)
@@
where $Q=(Q_{c}^{-1}+Q_{i}^{-1})^{-1}$ is the total quality factor of the resonator, which combines the coupling quality factor $Q_{c}$ and the internal quality factor $Q_{i}$ Zmuidzinas (2012). Finally, the parameter $y=Q\cdot\Delta f_{r}/f_{r}$ is the detuning of the readout frequency relative to the resonance width Zmuidzinas (2012); Mazin (2005); Vignati et al. (2021). Usually, the phase readout $\varphi$ provides a better signal-to-noise ratio Zmuidzinas (2012); therefore, this will be adopted in this work.
The signal phase responsivity $r$ of a KID is Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008); Delicato et al. (2025):
$r=\frac{\eta}{V_{\mathrm{KID}}}\frac{\alpha S_{2}(f_{r},T)Q}{N_{0}\Delta_{0}^{2}}\frac{1}{1+4y^{2}}$ (2)
@@
where $\eta$ is the energy to pair-breaking conversion efficiency, $V_{\mathrm{KID}}$ is the active inductor volume, $\alpha=L_{k}/(L_{k}+L_{g})$ is the kinetic inductance fraction, and $S_{2}(f_{r},T)$ is a dimensionless factor of order one from Mattis-Bardeen theory. Finally, $N_{0}$ is the single-spin density of states and $\Delta_{0}\simeq 1.76k_{B}T_{c}$ is the superconducting gap, with $k_{B}$ the Boltzmann constant and $T_{c}$ the critical temperature. We emphasize that in Eq. 2 the term $\Delta_{0}$ appears quadratically as the product of two identical gap factors: one term arises from the number of CPs broken by a given energy deposition $E$, scaling as $E/\Delta_{0}$, while the second originates from the single–spin volume density, $N_{0}\Delta_{0}$, within the sensor Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008).
@@
If phonons absorption is not saturated and losses due to supports or interfaces do not dominate, the efficiency $\eta$ scales almost linearly with the phonon collection volume $V_{\mathrm{ph}}$ Martinez et al. (2019); Fyhrie et al. (2016); Santis (2022):
@@
In a standard phonon-mediated KID, where $V_{\mathrm{ph}}$ coincides with the active volume $V_{\mathrm{KID}}$, the responsivity weakly depends on the inductor volume, as follows from Eq. 2. Conversely, separating these volumes allows to design devices with large $V_{\mathrm{ph}}$ and small $V_{\mathrm{KID}}$, which would enhance the detector responsivity.
@@
We developed and tested a phonon-mediated KID with separate phonon collectors Santis (2022); Mazin et al. (2006); Pesce et al. (2026) that act also as QPs funnels (see Fig. 1). We refer to this device as the FunKID and to the phonon collectors as the funnels. The resonator is composed of a 77 nm trilayer meander of Aluminum-Titanium-Aluminum (AlTiAl), with 14 nm-33 nm-30 nm thicknesses, respectively Cardani et al. (2018). The inductive part of the KID is composed of seven fingers of around $3600\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}$, evenly spaced by $215\,\mathrm{\mu m}$ (Fig. 1(a)). The funnels are Aluminum structures with $210\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}\times 100\,\mathrm{nm}$ volume intersecting the inductor, evenly spaced by $10\,\mathrm{\mu m}$ (Fig. 1(b)). The total phonon collection volume, which coincides with the total volume of the funnels, is $V_{\mathrm{ph}}\simeq 5.2\,\mathrm{mm}^{2}\times 100\,\mathrm{nm}$, while the active volume of the sensor is $V_{\mathrm{KID}}\simeq 0.55\,\mathrm{mm}^{2}\times 77\,\mathrm{nm}$, giving a volume ratio of $V_{\mathrm{ph}}/V_{\mathrm{KID}}\simeq 12$.
@@
Figure 1: (a) Layout of the FunKID. The funnels in purple intersect the inductive meander indicated with L, while the capacitor is indicated with C. (b) Detail of the funnels. The measures are in $\mathrm{\mu m}$.
@@
Due to proximity effects Cardani et al. (2018); Catalano, A. et al. (2015); Zhao et al. (2017); Brammertz et al. (2001) among the two Aluminum layers ($T_{c}\simeq 1.2\,\mathrm{K}$ Zhao et al. (2017)) and the central Titanium layer ($T_{c}\simeq 0.4\,\mathrm{K}$ Zhao et al. (2017)), the AlTiAl sensor has an intermediate critical temperature $T_{c}$, which is lower than that of the funnels, made of Aluminum. For the same design thicknesses, the CALDER project measured the critical temperature of an AlTiAl trilayer as $T_{c}\simeq 0.81\,\mathrm{K}$ Cardani et al. (2018, 2021); a similar value is expected for this resonator, assuming comparable boundary effects Zhao et al. (2018).
@@
In the FunKID design, schematically reported Fig. 2(a), the phonons produced in the substrate are mostly absorbed by the funnels, where they can break CP generating QPs, which, through diffusion, reach the sensor. Due to the energy gap difference between funnels with gap $\Delta_{\mathrm{Al}}$ and sensor with gap $\Delta_{\mathrm{AlTiAl}}$, QPs with energy $\Delta_{\mathrm{Al}}$ relax down to $\Delta_{\mathrm{AlTiAl}}$ and are trapped within the sensor, preventing their recombination in the funnels Mazin et al. (2006); Riwar and Catelani (2019).
@@
The length of the funnels was chosen to be smaller than the characteristic QP diffusion length in aluminum, $\lambda_{r}=\sqrt{D\tau_{r}}$ Nussbaumer et al. (2000), where $D$ is the QP diffusion coefficient and $\tau_{r}$ the QP recombination time. We adopted $D=22.5\,\mathrm{cm^{2}/s}$ Hsieh and Levine (1968) and $\tau_{r}\simeq 1\,\mathrm{ms}$ de Visser et al. (2011); Barends et al. (2008), yielding to a diffusion length of $\lambda_{r}\simeq 1.5\,\mathrm{mm}$, much higher than half the length of the funnels of $100\,\mathrm{\mu m}$. We can also define a characteristic time scale quantifying the maximum diffusion time of QPs inside the funnels before reaching the sensor, $\tau_{\mathrm{diff}}\sim l^{2}/D=4\,\mathrm{\mu s}$, where $l=100\,\mathrm{\mu m}$ is half of the funnel length. This sets the minimum timescale over which the sensor needs to integrate the signal.
@@
Figure 2: (a) Detection scheme of the FunKID. The white dots represent the athermal phonons generated by an energy deposition in the substrate, the red paired dots are the CPs and the single red dots the QPs diffusing from the funnel to the KID. More details in the text. (b) Current density simulated with SONNET. The length reported is measured in $\mathrm{\mu m}$.
@@
The responsivity gain of the FunKID relative to an identical standard KID without funnels (with phonon collection volume $V_{\mathrm{KID}}$), $r_{\mathrm{FunKID}}/r_{\mathrm{KID}}$, is found using Eq. 2:
$\frac{r_{\mathrm{FunKID}}}{r_{\mathrm{KID}}}\simeq\frac{V_{\mathrm{ph}}}{V_{\mathrm{KID}}}\frac{\alpha_{\mathrm{FunKID}}}{\alpha_{\mathrm{KID}}}\frac{Q_{\mathrm{FunKID}}}{Q_{\mathrm{KID}}}\frac{\Delta_{\mathrm{AlTiAl}}}{\Delta_{\mathrm{Al}}}$ (4)
where we assumed that $N_{0}$ and $S_{2}(f_{r},T)$ were approximately equal between the two devices and that their response was linear, thus $1+4y^{2}\simeq 1$. Furthermore, we considered only the geometric dependence of $\eta$ for both KIDs, according to Eq. 3. Under identical electrical conditions, indeed equal quality factors $Q$ and kinetic inductance fractions $\alpha$ for both devices we expect a responsivity gain of approximately 8.
@@
The test device was fabricated at the PTA facility clean room in Grenoble Plateforme technologique amont using a $500\,\mathrm{\mu m}$ thick silicon tile. On the same substrate, an identical standard phonon-mediated KID, with a simulated resonant frequency $20\,\mathrm{MHz}$ higher than the FunKID, was also produced (Fig. 3). The KID sensors and the feed-line were etched from the 77 nm AlTiAl trilayer following a standard lithographic technique. After their fabrication, photo-resist was used to coat the surface of the wafer, and gaps in the shape of the funnels were patterned on it. The funnels were then realized by selectively depositing $100\,\mathrm{nm}$ of Aluminum in the patterned gaps, following a lift-off process.
The tile was mounted into a copper holder by means of Teflon supports and wire bonded to $50\,\mathrm{\Omega}$ launchers. The device was operated in a dilution $^{3}\mathrm{He}/^{4}\mathrm{He}$ cryostat, at a base temperature $T_{0}\simeq 35\,\mathrm{mK}$. Finally, a mu-metal shield was placed outside the cryostat to reduce magnetic-field effects Flanigan et al. (2016).
To calibrate the detectors, three optical fibers were placed facing the substrate backside: two were aligned with the resonators for their identification Pesce et al. (2026) and one was located at the center to calibrate the KIDs from the same position for comparison. The fibers were pulsed with $400\,\mathrm{nm}$ photons bursts generated at room temperature by a controlled LED lamp Del Castello (2024).
The readout tones were generated by the NIXA superheterodyne radio-frequency detection system Bourrion et al. (2011, 2013) transmitting and receiving tones at the KIDs resonant frequencies. The transmitted waves undergo an attenuation chain of $55\,\mathrm{dB}$ before the chip, are transmitted past the resonators, then are amplified by 30 dB by a Low Noise Amplifier Cosmic Microwave Technology, Inc. (2025) operated at 4 K and finally received back by NIXA.
The resonances of the two KIDs are shown in Fig. 4. We fitted both the amplitude of the transmission function $|S_{21}|$ and its real and imaginary components in order to extract the resonant frequencies and the quality factors $Q_{c}$, $Q_{i}$, and $Q$ of the two KIDs. We also evaluated their resonator ring times $\tau_{\mathrm{ring}}=Q/\pi f_{r}$. The results are reported in Tab. 1.
@@
Table 1: Results of the resonant frequency $f_{r}$, quality factors $Q$, $Q_{i}$ and $Q_{c}$ and ring times $\tau_{\mathrm{ring}}$ of the two resonators.
A shift of around $40\,\mathrm{MHz}$ of the measured FunKID resonant frequency with respect to the simulation was observed, indicating that the overlap regions between the funnels and the resonator might have not been correctly modeled in the simulations. Conversely, the obtained KID resonant frequency is in agreement with simulation, assuming a kinetic inductance $L_{k}=1.4\,\mathrm{pH/sq}$, which is the same of Ref. Cardani et al. (2018). The coupling quality factors $Q_{c}\sim 20\,\mathrm{k}$ are in line with the values simulated with SONNET, for both resonators. The internal quality factor of the FunKID is comparable to the coupling quality factor, which currently limits this design. Conversely, the internal quality of the KID is higher than the FunKID and in line with the values obtained in Ref. Cardani et al. (2018), indicating that the reduced $Q_{i}$ of the FunKID is set by the funnels and not by fabrication processes of the trilayer AlTiAl film.
For each resonator, we evaluated an estimate of kinetic inductance fraction $\alpha$, by combining the measured resonant frequencies $f_{r}$ with the resonant frequencies $f_{0}$ obtained from SONNET simulations performed with a perfect conductor ($L_{k}=0$Cardani et al. (2018):